# Optical next generation reservoir computing

### v2

@@ Results
Figure 4: Optical NGRC observer. a For a dynamical system, often partial information of the full state of the system is measurable, e.g., state variables $[u_{1},...,u_{k}]^{T}$ are observables while $[u_{k+1},...,u_{M}]^{T}$ are unmeasured. The optical NGRC extracts information from measured observables (blue) and predicts unmeasured variables (purple) based on the state of the reservoir (orange). b Two variables $u_{1}$ and $u_{2}$ (blue) of the Lorenz63 system are provided as observables to infer the third variable $u_{3}$. The predicted output by optical NGRC observer (red) matches the ground truth (blue) with high accuracy (NRMSE $=$ 0.0169). c The optical NGRC observer results of the KS time series. 7 out of 64 spatial grids (evenly spaced in the spatial dimension) are input of the optical NGRC to infer the remaining 57 unmeasured variables. Top: ground truth; Middle: reservoir prediction (also including the observables for clarity); Bottom: error. d Performance comparison of the optical NGRC observer and the spline interpolation on the KS time series. At different number of observables, the Pearson correlation between the optical NGRC observer and the grouth truth is consistently higher than that of the spline interpolation and the ground truth.
@@
Regarding its long-term prediction performance, we illustrate in Fig. 3b a section (spanning from $t_{1}$ to $t_{2}$) of predicted outputs for 10,000 time steps (starting at $t_{0}$) beyond the short-term regime. While the prediction completely deviates from the ground truth at the element-wise level, the visual inspection indicates that the optical NGRC captures the correct ‘climate’ [72]. We substantiate this observation by quantitatively analyzing the power spectra of the predicted outputs, the KS ground truth and a random noise signal in Fig. 3c (see data processing details in Methods). The long-term prediction results presented in Figs. 2 and 3 indicate that the optical NGRC effectively synchronizes with host prototypical systems, functioning as a physical twin without knowing their models.
@@
Optical NGRC observer. We now proceed to the application of optical NGRC in a third benchmark task, referred to as the ‘reservoir observer’ [17, 70]. As illustrated in Fig. 4a, in many contexts when studying a dynamical system, it is common to have access to only a partial set of its complete degrees of freedom at a given time. An ‘observer’ aims to deduce unmeasured variables from the measured ones (i.e., observables), for example, $[u_{1},...,u_{k}]^{T}\xrightarrow{\mathrm{optical\thinspace NGRC}}[u_{k+1},...,u_{M}]^{T}$ (see Supplementary Algorithm 2). As before, we first train the optical reservoir in a supervised fashion, based on the limited number of time measurements where the full state variables of the system $[u_{1},...,u_{k},...,u_{M}]^{T}$ are accessible. Here, we conduct optical NGRC observer experiments on both the Lorenz63 and KS systems.
@@
To follow the convention in digital NGRC [70], we employ one current and four delayed inputs, uniformly sampled with a stride of five time steps. For the Lorenz63 system, we infer $u_{3}$ from $u_{1}$ and $u_{2}$. Figure 4b shows that only a short period of training time with 400 time steps yields decent predictions for $6020<t<$ 60, manifesting the feasibility of our optical NGRC in this application. Going beyond, we investigate the optical NGRC observer in the spatiotemporal KS system, based on the sparse spatial information that is available. Specifically, for the domain size of $L=22$ of the KS time series studied in this work, we sample $64$ spatial points at each time step. The experiments are performed using the knowledge of uniformly sampled $S$ spatial state variables to infer the remaining set of $64-S$ variables. Figure 4c presents the test results when $S=7$. We also vary $S$ from 4 to 8 and summarize the calculated correlation between the experimental results and the corresponding ground truth in Fig. 4d. To provide a digital baseline, the cubic spline interpolation method is also implemented (see data processing and performance metrics details in Methods). We observe that the optical NGRC observer consistently outperforms the spline interpolation, serving as an effective means to reconstruct unmeasured dynamical system variables.
@@ Discussion
Generally, for physical RC systems, experimental factors such as device quantization and noise can limit the RC performance. Nevertheless, these requirements are not overly stringent for optical NGRC. In Supplementary Note 4, we show that 7- or 8-bit depth SLM and camera are sufficient for the tasks demonstrated in this work. To reduce the effect of the noise, we optimize the setup and average the experimental measurements when deriving the reservoir states (see Supplementary Note 2). Beyond forecasting chaotic time series, it is interesting to explore other machine learning tasks with optical NGRC, such as speech recognition [75] and graph classification [69, 65]. At the architecture level, the optical NGRC can also be chained to create more advanced neural networks like deep or parallel NGRC [69, 76] for enhanced expressivity and versatile functionalities [77].
@@
Compared to digital NGRC, the scaling of optical NGRC and consequently the computational complexity, exhibits notable distinctions. Specifically, digital NGRC features polynomial scaling with respect to the input data dimension (see Supplementary Note 5), provided that dimension reduction of input data is not performed based on any prior knowledge [71, 70]. This imposes a progressively heavier burden for digital NGRC to process large-scale systems. For example, for the same KS time series prediction studied in this work, the original digital NGRC scheme would require a reservoir size of 8,385 when considering up to second-order polynomial terms, while we only use 2,500 in optical NGRC (see Supplementary Note 5). More interestingly, due to the use of phase encoding and intensity detection in the optical NGRC, higher-order polynomial terms of inputs (beyond the second-order) are naturally embedded in every speckle grain mode, readily accessible by the linear readout layer. As such, we do not have to manually determine the needed polynomial order or select the terms as necessitated in digital NGRC. In addition, our optical NGRC setup can be readily scaled to large input and output dimensions, where we can leverage the potential of optical computing with scattering media as discussed in refs. [78, 63, 79].
@@
In summary, we propose and experimentally demonstrate in this work an efficient optical NGRC scheme based on light scattering through disordered media. Similar to the spirit of digital NGRC, optical NGRC generates the mixture of polynomial functions of time-delayed inputs. We leverage these features embedded in optical speckles for various benchmark tasks in RC, ranging from short-term and long-term forecasting to reservoir observer in Lorenz63 and KS chaotic systems. Optical NGRC features several advantages over conventional optical RC in significantly shorter training length, fewer hyperparameters, increased interpretability and greater expressivity, and shows the prospect in scalability compared to digital NGRC towards large-scale chaotic systems (see Supplementary Table S1). Broadly, the proposed optical NGRC framework is hardware-agnostic, opening up new possibilities for a wide variety of physical computing substrates.

### v3

@@ Results
Figure 4: Optical NGRC observer. (a) For a dynamical system, often partial information of the full state of the system is measurable, e.g., state variables $[u_{1},...,u_{k}]^{T}$ are observables while $[u_{k+1},...,u_{M}]^{T}$ are unmeasured. The optical NGRC extracts information from measured observables (blue) and predicts unmeasured variables (purple) based on the state of the reservoir (orange). (b) Two variables $u_{1}$ and $u_{2}$ (blue) of the Lorenz63 system are provided as observables to infer the third variable $u_{3}$. The predicted output by optical NGRC observer (red) matches the ground truth (blue) with high accuracy (NRMSE $=$ 0.0169). (c) The optical NGRC observer results of the KS time series. 7 out of 64 spatial grids (evenly spaced in the spatial dimension) are input of the optical NGRC to infer the remaining 57 unmeasured variables. Top: ground truth; Middle: reservoir prediction (also including the observables for clarity); Bottom: error. (d) Performance comparison of the optical NGRC observer and the spline interpolation on the KS time series. The Pearson correlation between the optical NGRC observer prediction and the ground truth is consistently higher than that between the spline interpolation and the ground truth.
@@
Regarding its long-term prediction performance, we illustrate in Fig. 3(b) a section (spanning from $t_{1}$ to $t_{2}$) of predicted outputs for 10,000 time steps (starting at $t_{0}$) beyond the short-term regime. While the prediction completely deviates from the ground truth at the element-wise level, the visual inspection indicates that the optical NGRC captures the correct ‘climate’ [57]. We substantiate this observation by quantitatively analyzing the power spectra of the predicted outputs, the KS ground truth and a random noise signal in Fig. 3(c) (see data processing details in Methods). The long-term prediction results presented in Figs. 2 and 3 indicate that the optical NGRC effectively synchronizes with host prototypical systems, functioning as a physical twin without knowing their models.
@@
Optical NGRC observer. We now proceed to the application of optical NGRC in a third benchmark task, referred to as the ‘reservoir observer’ [14, 53]. As illustrated in Fig. 4(a), in many contexts when studying a dynamical system, it is common to have access to only a partial set of its complete degrees of freedom at a given time. An ‘observer’ aims to deduce unmeasured variables from the measured ones (i.e., observables), for example, $[u_{1},...,u_{k}]^{T}\xrightarrow{\mathrm{optical\thinspace NGRC}}[u_{k+1},...,u_{M}]^{T}$ (see Supplementary Algorithm 2). As before, we first train the optical reservoir in a supervised fashion, based on the limited number of time measurements where the full state variables of the system $[u_{1},...,u_{k},...,u_{M}]^{T}$ are accessible. Here, we conduct optical NGRC observer experiments on both the Lorenz63 and KS systems.
@@
To follow the convention of the observer task in digital NGRC [53], we employ one current and three delayed inputs to infer unmeasured variables, rather than the two inputs used in previous autonomous forecasting tasks. These four inputs are uniformly sampled with a stride of five time steps (see details in Supplementary Algorithm 2). For the Lorenz63 system, we infer $u_{3}$ from $u_{1}$ and $u_{2}$. Figure 4(b) shows that only a short period of training time with 400 time steps yields decent predictions for $6020<t<$ 60, manifesting the feasibility of our optical NGRC in this application. Going beyond, we investigate the optical NGRC observer in the spatiotemporal KS system, based on the sparse spatial information that is available. Specifically, for the domain size of $L=22$ of the KS time series studied in this work, we sample $64$ spatial points at each time step. The experiments are performed using the knowledge of uniformly sampled $S$ spatial state variables to infer the remaining set of $64-S$ variables. Figure 4(c) presents the test results when $S=7$. We also vary $S$ from 4 to 8 and summarize the calculated correlation between the experimental results and the corresponding ground truth in Fig. 4(d). To provide a digital baseline, the cubic spline interpolation method is also implemented (see data processing and performance metrics details in Methods). We observe that the optical NGRC observer consistently outperforms the spline interpolation, serving as an effective means to reconstruct unmeasured dynamical system variables.
@@ Discussion
Generally, for physical RC systems, experimental factors such as device quantization and noise can limit the RC performance. Nevertheless, these requirements are not overly stringent for optical NGRC. In Supplementary Note 4, we show that 7- or 8-bit depth SLM and camera are sufficient for the tasks demonstrated in this work. To reduce the effect of the noise, we optimize the setup and average the experimental measurements when deriving the reservoir states (see Supplementary Note 2). Beyond forecasting chaotic time series, it is interesting to explore other machine learning tasks with optical NGRC, such as speech recognition [60] and graph classification [52, 48]. At the architecture level, the optical NGRC can also be chained to create more advanced neural networks like deep or parallel NGRC [52, 61] for enhanced expressivity and versatile functionalities [62]. For instance, we can employ a thin scattering medium to only introduce the random mixing locally [63], so as to implement parallel optical NGRC [61].
It is important to note that many steps in our current implementation involve processing in a digital computer like many other optical RC systems. Upgrading the system to an all-optical version remains a promising challenge, which could further improve the energy efficiency and computing throughput. Currently, we employ only a small fraction of pixels of the encoding SLM and detection camera. Our optical NGRC system can be scaled to accommodate larger inputs and larger reservoir output dimensions, allowing us to explore the potential scaling advantages of optical computing with scattering media as discussed in refs. [64, 46, 65]. While edge computing hardware have shown excellent performance in implementing NGRC [66], it is restricted to accelerating relatively small-scale tasks and optical systems comprising more advanced components may offer further improvements (see Supplementary Note 5).
@@
Compared to digital NGRC, the scaling of optical NGRC and consequently the computational complexity, exhibits notable distinctions. Specifically, digital NGRC features polynomial scaling with respect to the input data dimension (see Supplementary Note 5), provided that dimension reduction of input data is not performed based on any prior knowledge [54, 53]. This could impose a progressively heavier burden for digital NGRC to process large-scale systems. For example, for the same KS time series prediction studied in this work, the original digital NGRC scheme would require a reservoir size of 8,385 when considering up to second-order polynomial terms, while we only use 2,500 in optical NGRC (see Supplementary Note 5). More interestingly, due to the use of phase encoding and intensity detection in the optical NGRC, higher-order polynomials functions of delayed inputs (beyond the second-order and in principle infinite orders due to Taylor expansion) are naturally embedded in every speckle grain mode, and they can be readily accessed by the linear readout layer if needed. As such, we do not have to manually determine the needed polynomial order or select the terms as necessitated in digital NGRC. In our case, we may use the phase encoding range as another hyperparameter to optimize the reservoir.
@@
In summary, we propose and experimentally demonstrate in this work an efficient optical NGRC scheme based on light scattering through disordered media. Similar to the spirit of digital NGRC, optical NGRC generates the mixture of polynomial functions of time-delayed inputs. We leverage these features embedded in optical speckles for various benchmark tasks in RC, ranging from short-term and long-term forecasting to reservoir observer in Lorenz63 and KS chaotic systems. Optical NGRC features several advantages over conventional optical RC based on scattering media in significantly shorter training length, fewer hyperparameters, increased interpretability and greater expressivity, and may hold the prospect in scalability compared to digital NGRC towards large-scale chaotic systems (see Supplementary Table S1). Broadly, the proposed optical NGRC framework is hardware-agnostic, which could inspire new possibilities for a wide variety of physical RC substrates.